# Balancing Embedding Spectrum for Recommendation

### v1

@@ IV-B A Decorrelation Perspective
$\mathbf{H}_{U}\leftarrow\mathbf{H}_{U}-\alpha\left(\mathbf{H}_{U}\mathbf{H}_{U}^{T}\right)\mathbf{H}_{U},$ (16)
@@
$\mathbf{H}_{I}\leftarrow\mathbf{H}_{I}-\alpha\left(\mathbf{H}_{I}\mathbf{H}_{I}^{T}\right)\mathbf{H}_{I},$
where $\mathbf{H}_{U}$ and $\mathbf{H}_{I}$ are user and item embedding matrices, respectively. Therefore, Equation (14) is equivalent to DirectSpec. Intuitively, two users/items are (positively or negatively) correlated if $|\mathbf{h}_{u}^{T}\mathbf{h}_{v}|>0$ or $|\mathbf{h}_{i}^{T}\mathbf{h}_{j}|>0$, then we can see that the goal of Equation (14) is to decorrelate/orthogonalize different users and items (e.g., $\mathbf{h}_{u}^{T}\mathbf{h}_{v}\rightarrow 0$ and $\mathbf{h}_{i}^{T}\mathbf{h}_{j}\rightarrow 0$) instead of simply pushing them away.
Proposition 4.
The lower bound of $L_{ds}\rightarrow 0$ as $d\rightarrow{\rm\max}(|\mathcal{U}|,|\mathcal{I}|)$.
We can see that different users/items are orthogonal to each other when $L_{ds}=0$. Then, consider the following linear combination:
$\alpha_{1}\mathbf{h}_{1}+\cdots+\alpha_{n}\mathbf{h}_{n}=0.$ (17)
By multiplying Equation (17) by $\mathbf{h}_{k}^{T}$, we get $\alpha_{k}=0$, indicating that orthogonal vectors are linearly independent as well, requiring $d\geq{\rm\max}(|\mathcal{U}|,|\mathcal{I}|)$.
From a different perspective, $\mathbf{H}_{U}\mathbf{H}_{U}^{T}$ and $\mathbf{H}_{I}\mathbf{H}_{I}^{T}$ represent the similarity scores between users and items, respectively, and we have the following relations according to SVD:
$\mathbf{H}_{U}\mathbf{H}_{U}^{T}=\mathbf{P}diag(\sigma_{k}^{2})\mathbf{P}^{T},\,\mathbf{H}_{I}\mathbf{H}_{I}^{T}=\mathbf{Q}diag(\sigma_{k}^{2})\mathbf{Q}^{T}.$ (18)
For simplicity, we use the same notations denoting singular values and vectors of $\mathbf{H}_{U}$ and $\mathbf{H}_{I}$, since we do not emphasize their difference here. In the meanwhile, $\mathbf{P}$ and $\mathbf{Q}$ are also the eigenvectors of $\mathbf{H}_{U}$ and $\mathbf{H}_{I}$, respectively. Let $\lambda_{k}$ denote the eigenvalue of $\mathbf{H}_{U}\mathbf{H}_{U}^{T}$ and $\mathbf{H}_{I}\mathbf{H}_{I}^{T}$, we have $\lambda_{k}=\sigma_{k}^{2}$. When $(\mathbf{H}_{U}\mathbf{H}_{U}^{T})_{uv}\rightarrow 0$ and $(\mathbf{H}_{I}\mathbf{H}_{I}^{T})_{ij}\rightarrow 0$, $(\mathbf{H}_{U}\mathbf{H}_{U}^{T})$ and $(\mathbf{H}_{I}\mathbf{H}_{I}^{T})$ are optimized to be identity matrices with uniform spectrum distributions: $\lambda_{1}=\lambda_{2}=\cdots=1$, thus $\mathbf{H}_{U}$’s and $\mathbf{H}_{I}$’s spectra are also uniform, leading to a maximum erank. The above observations reveal that unobserved pairs should stay irrelevant instead of simply being pushed away like BCE and BPR loss, which is more consistent with users’ true preference over unobserved interactions. We will empirically compare DirectSpec and negative sampling in Section V-B3.
IV-C DirectSpec+
As the observed pairs should be deeply correlated, perfect user/item decorrelation is not the goal of personalized recommendation, explaining why we do not need (1) the erank of the embedding to strictly be $d$ and (2) $d$ to be large enough to reach the perfect decorrelation. The training objective should balance between user/item correlation and decorrelation:
$L=\sum_{(u,i)\in\mathbf{R}^{+}}\ln\sigma\left(\mathbf{h}_{u}^{T}\mathbf{h}_{i}\right)+L_{ds}.$ (19)
The first term pulls users/items to be close to each other, while the second term offsets the negative effect from the first term by orthogonalizing them. From Equation (15), we can see that all samples are optimized with the same pace. To enhance the efficiency, we can employ self-paced gradients according to the extent of correlation (e.g., $\textbf{h}_{u}^{T}\textbf{h}_{v}\rightarrow 1$ and $\textbf{h}_{i}^{T}\textbf{h}_{j}\rightarrow 1$ should be penalized more) by improving Equation (14) as:
$L_{ds}=\sum_{u\in\mathcal{U}}\sum_{v\in\mathcal{U}}\exp\left(\mathbf{h}_{u}^{T}\mathbf{h}_{v}/\tau\right)+\sum_{i\in\mathcal{I}}\sum_{j\in\mathcal{I}}\exp\left(\mathbf{h}_{i}^{T}\mathbf{h}_{j}/\tau\right).$ (20)
Compared with Equation (15), the pairs with $\textbf{h}_{u}^{T}\textbf{h}_{v}\rightarrow 1$ and $\textbf{h}_{i}^{T}\textbf{h}_{j}\rightarrow 1$ are pushed away at a faster pace (i.e., larger gradients); $\tau$ is a temperature controlling the strength of penalties. We focus on penalizing the positive correlation as optimization on the observed interactions causes user/item pairs to be positively correlated. Note that Equation (20) is equivalent to a pairwise Gaussian potential: $\exp\left(\frac{\mathbf{h}_{u}^{T}\mathbf{h}_{v}}{\tau}\right)=\exp\left(\frac{-\|\mathbf{h}_{u}-\mathbf{h}_{v}\|^{2}}{2\tau}+\frac{1}{\tau}\right)$, where the minimum is reached when user/item embeddings are orthogonal to each other when $d$ is large enough according to Proposition 3 in [20]. Thus, the training objective of the enhanced algorithm is consistent with the original DirectSpec. Furthermore, we notice that some unobserved pairs are correlated as well to some extent, such as the users showing similar preference or items interacted by similar users that should be decorrelated with a slower pace than other irrelevant pairs. Here, we propose two adaptive temperature designs to adapt to the pairs with different degrees of correlations:
@@ IV-C DirectSpec+
    We define the graph distance between two nodes: $d_{\mathcal{G}}\left(\cdot,\cdot\right)$ as the minimum length of path between them on $\mathcal{G}$, then an unparameterrized design is defined as follows:
    $\mathbf{\Gamma}_{uv}=\left\{\begin{array}[]{lr}\tau_{0}\quad\quad\,\,d_{\mathcal{G}}\left(u,v\right)\leq K\\ \\ \tau_{1}\quad\quad\ otherwise,\\ \end{array}\right.$ (21)
    where the pairs with $d_{\mathcal{G}}\left(u,v\right)\leq K$ are considered correlated, and $\tau_{0}\leq\tau_{1}$.
By calculating the gradients of Equation (20), we can incorporate its parameter-updating formula into Algorithm 2, and propose DirectSpec+:
$\mathbf{H}_{B}^{U}\leftarrow\mathbf{H}_{B}^{U}-\alpha\cdot{\rm softmax}\left(\mathbf{H}_{B}^{U}{\mathbf{H}_{B}^{U}}^{T}\odot\mathbf{\Gamma}\right)\mathbf{H}_{B}^{U},$ (22)
$\mathbf{H}_{B}^{I}\leftarrow\mathbf{H}_{B}^{I}-\alpha\cdot{\rm softmax}\left(\mathbf{H}_{B}^{I}{\mathbf{H}_{B}^{I}}^{T}\odot\mathbf{\Gamma}\right)\mathbf{H}_{B}^{I}.$
To regulate the magnitude of the gradient, we rewrite it as a Softmax function. Other steps of DirectSpec+ are consistent with Algorithm 2.
IV-D Discussion
IV-D1 Connection and Comparison Between DirectSpec+ and Uniformity Loss
Contrastive learning (CL) has shown great potential in recommender systems [21, 22], which adopts an InfoNCE loss:
$L_{cl}=-\sum_{u}\ln\frac{\exp\left(\mathbf{h}_{u}^{T}\mathbf{h}_{u^{+}}/\tau\right)}{\sum_{v}\exp\left(\mathbf{h}_{u}^{T}\mathbf{h}_{v}/\tau\right)}-\sum_{i}\ln\frac{\exp\left(\mathbf{h}_{i}^{T}\mathbf{h}_{i^{+}}/\tau\right)}{\sum_{j}\exp\left(\mathbf{h}_{i}^{T}\mathbf{h}_{j}/\tau\right)},$ (23)
that can be decoupled to alignment and uniformity loss. Alignment minimizes the distance between positive pairs (i.e., $(u,u^{+})$ and $(i,i^{+})$). In the meanwhile, embeddings from different users/items are pushed away via uniformity loss, which has been shown necessary for desirable user/item representations [23, 24]. To unveil the effectiveness of uniformity, we fix alignment: $\mathbf{h}_{u}^{T}\mathbf{h}_{u^{+}}=1$ and $\mathbf{h}_{i}^{T}\mathbf{h}_{i^{+}}=1$, and let $L_{cl}(u)=-\ln\frac{\exp\left(1/\tau\right)}{\sum_{v}\exp\left(\mathbf{h}_{u}^{T}\mathbf{h}_{v}/\tau\right)}$. Then, the gradients over the parameters are calculated as follows:
$\frac{\partial L_{cl}}{\partial\mathbf{h}_{u}}=\frac{\partial L_{cl}(u)}{\partial\mathbf{h}_{u}}+\sum_{v\neq u}\frac{\partial L_{cl}(v)}{\partial\mathbf{h}_{u}}$ (24)
$=\sum_{v}\frac{\exp\left(\mathbf{h}_{u}^{T}\mathbf{h}_{v}/\tau\right)}{\sum_{k}\exp\left(\mathbf{h}_{u}^{T}\mathbf{h}_{k}/\tau\right)}\mathbf{h}_{v}/\tau+\sum_{v}\frac{\exp\left(\mathbf{h}_{u}^{T}\mathbf{h}_{v}/\tau\right)}{\sum_{k}\exp\left(\mathbf{h}_{v}^{T}\mathbf{h}_{k}/\tau\right)}\mathbf{h}_{v}/\tau,$
thus the parameters are updated via SGD as:
$\mathbf{H}_{U}\leftarrow\mathbf{H}_{U}-\alpha\cdot\left({\rm softmax}\left(\mathbf{H}_{U}{\mathbf{H}_{U}}^{T}/\tau\right)+\mathbf{S}\mathbf{D}^{-1}\right)\mathbf{H}_{U},$ (25)
where $\mathbf{S}=\exp\left(\mathbf{H}_{U}{\mathbf{H}_{U}}^{T}/\tau\right)$, and $\mathbf{D}$ is a diagonal matrix with $D_{uu}=\sum_{v}\mathbf{S}_{uv}$, ${\rm softmax}\left(\mathbf{H}_{U}{\mathbf{H}_{U}}^{T}/\tau\right)=\mathbf{D}^{-1}\mathbf{S}$. Item embeddings are updated in a way similar to Equation (25). We can see that Equation (25) is the same as DirectSpec+ if we ignore the term $\mathbf{S}\mathbf{D}^{-1}$, revealing that uniformity can alleviate collapse by indirectly rebalancing the embedding spectrum.
The superiority of DirectSpec over CL-based methods are two aspects. Firstly, CL-based methods generate and compare different node views based on data augmentation, requiring more complexity than our methods. In addition, DirectSpec is more effective tackling the embedding collapse issue as we showed that CL-based methods rebalance the embedding spectrum in a more indirect and implicit way. We will demonstrate the above two claims in Section V-B.
TABLE II: Comparison of time complexity.
Complexity/Model DirectSpec CL GCL
Self-supervised $c|\mathbf{R}^{+}|Bd$ $c|\mathbf{R}^{+}|\left(|\mathcal{U}|+|\mathcal{I}|\right)d$ or $c|\mathbf{R}^{+}|Bd$ $c|\mathbf{R}^{+}|\left(\frac{L|\mathbf{R}^{+}|}{B}\right.$ $\left.+|\mathcal{U}|^{2}+|\mathcal{I}|^{2}\right)d$
Supervised $c|\mathbf{R}^{+}|d$ $2c|\mathbf{R}^{+}|d$ or $c|\mathbf{R}^{+}|Bd$ $2c|\mathbf{R}^{+}|d$
IV-D2 Complexity
We compare the complexity of DirectSpec with graph contrastive learning (GCL) and CL in Table II, where $B$ and $c$ denote the batch size and the required epochs, respectively. All the three methods adopt a multi-task training strategy, where the complexity comes from the supervised and self-supervised learning task (if we consider the process of spectrum balancing in DirectSpec as a self-supervised learning task). For self-supervised learning task, GCL has expensive computational cost as it generates two node views on perturbed graphs, whose time complexity is twice that of the backbone GCN (here, we use LightGCN). CL has lower time complexity than GCL as it does not exploit graph structures, and the complexity can be reduced to that of DirectSpec: $c|\mathbf{R}^{+}|Bd$ if it adopts a mini-batch training [22] (the complexity in practical should be higher as we do not consider the complexity for data augmentation here). For supervised learning task, we ignore the computational cost of the encoder (i.e., the complexity for generating the embeddings) as DirectSpec and CL can also be implemented on other algorithms (e.g., MLP, GNN, transformer, etc.). Compared with most of the works adopting BPR or CL loss which still rely on negative samples, DirectSpec only samples the positive pairs and is equipped with a simple loss. Overall, DirectSpec shows lower time complexity over GCL and CL methods.
IV-D3 Comparison between Embedding Collapse and Over-Smoothing
Embedding collapse is a general issue on recommendation algorithms that can be mathematically formulated as follows:
${\rm rank}\left(\mathbf{H}^{(L)}\right)\leq\epsilon,$ (26)
where $\mathbf{H}^{(L)}$ is the embedding when erank converges after updating $L$ times, $\epsilon<d$, and a smaller $\epsilon$ indicates a more serious collapse. On the other hand, over-smoothing only exists on GNN-based methods caused by repeating the message passing rule, which can be considered as a complete collapse issue and be formulated as Equation (26) ($\epsilon\ll d$). Therefore, our proposed DirectSpec is expected to well address over-smoothing as well. Existing solutions alleviate the over-smoothing mostly by randomly dropping out the edges to perturb the spectrum [25] or balancing the contribution of different order neighborhood [26] to avoid all eigenvalues from shrinking. We can see that these methods can only slow the speed of collapse or prevent user/item representations from being the same (i.e., complete collapse), while there is no guarantee that incomplete collapse can also be prevented (i.e., $d>\tau>1$). We will demonstrate the effectiveness of our methods for over-smoothing in Section V-B3.
V Experiments
In this section, we comprehensively evaluate our proposed methods on three public datasets. We implement DirectSpec on two popular baselines: MF and LightGCN. Particularly, we introduce data description and implementation details in Section V-A; we compare DirectSpec with other competitive baselines in terms of accuracy and efficiency in Section V-B. Finally, we conduct model analysis in Section V-C, including detailed analysis and experimental results on how DirectSpec prevents embeddings from collapsing without relying on negative samples, and how different settings of hyperparameters affect model performance.
TABLE III: Statistics of datasets.
@@ V Experiments
TABLE IV: Performance comparison. The best baseline is underlined. “*” indicates statistical significance at $p<0.01$ for a one-tailed t-test.
CiteULike Yelp Gowalla
nDCG@$k$ Recall@$k$ nDCG@$k$ Recall@$k$ nDCG@$k$ Recall@$k$
$k$=10 $k$=20 $k$=10 $k$=20 $k$=10 $k$=20 $k$=10 $k$=20 $k$=10 $k$=20 $k$=10 $k$=20
BPR 0.1620 0.1773 0.1778 0.2190 0.0487 0.0583 0.0607 0.0869 0.1164 0.1255 0.1186 0.1483
CL-REC 0.1523 0.1662 0.1671 0.2069 0.0476 0.0566 0.0588 0.0833 0.1116 0.1198 0.1123 0.1401
CCL(MF) 0.1545 0.1642 0.1716 0.1996 0.0509 0.0593 0.0617 0.0846 0.1269 0.1349 0.1295 0.1573
EMNF-U 0.1335 0.1601 0.1556 0.2193 0.0665 0.0811 0.0786 0.1167 0.1166 0.1286 0.1173 0.1528
DirectAU(MF) 0.2102 0.2252 0.2260 0.2693 0.0721 0.0848 0.0872 0.1219 0.1286 0.1402 0.1349 0.1710
DirectAU(GCN) 0.1926 0.2109 0.2116 0.2604 0.0695 0.0819 0.0854 0.1194 0.1298 0.1409 0.1363 0.1711
LightGCN 0.1610 0.1781 0.1771 0.2190 0.0572 0.0690 0.0721 0.1045 0.0987 0.1098 0.1074 0.1399
SGL-ED 0.1890 0.2065 0.2117 0.2588 0.0676 0.0794 0.0837 0.1166 0.1343 0.1462 0.1417 0.1779
LightGCL 0.2096 0.2238 0.2214 0.2638 0.0673 0.0790 0.0836 0.1151 0.1368 0.1482 0.1436 0.1801
GDE 0.1890 0.2061 0.2055 0.2528 0.0653 0.0771 0.0805 0.1129 0.1261 0.1367 0.1313 0.1656
LogDet 0.1403 0.1579 0.1548 0.2034 0.0711 0.0833 0.0863 0.1203 0.1179 0.1272 0.1203 0.1151
DirectSpec(MF) 0.1688 0.1849 0.1827 0.2270 0.0689 0.0804 0.0839 0.1156 0.1133 0.1225 0.1183 0.1480
DirectSpec(GCN) 0.1693 0.1863 0.1875 0.2334 0.0712 0.0832 0.0861 0.1192 0.1160 0.1251 0.1191 0.1490
DirectSpec+(MF) 0.2197 0.2354 0.2352 0.2818 0.0743 0.0864 0.0909 0.1249 0.1389 0.1509 0.1447 0.1829
DirectSpec+(GCN) 0.2038 0.2213 0.2213 0.2704 0.0745 0.0868 0.0909 0.1252 0.1400 0.1513 0.1453 0.1819
V-A Experimental Settings
V-A1 Datasets

### v2

@@ 4.3. Individualized Spectrum Balancing (DirectSpec+)
(20) $\mathbf{H}_{B}^{U}\leftarrow\mathbf{H}_{B}^{U}-\alpha\cdot{\rm softmax}\left(\mathbf{H}_{B}^{U}{\mathbf{H}_{B}^{U}}^{T}\odot\mathbf{\Gamma}\right)\mathbf{H}_{B}^{U},$
@@
$\mathbf{H}_{B}^{I}\leftarrow\mathbf{H}_{B}^{I}-\alpha\cdot{\rm softmax}\left(\mathbf{H}_{B}^{I}{\mathbf{H}_{B}^{I}}^{T}\odot\mathbf{\Gamma}\right)\mathbf{H}_{B}^{I},$
where $\odot$ represents the element-wise multiplication. Here, the individualized coefficient between $i$ and $j$ is replaced by $\frac{\exp(\tau_{ij}\cdot\mathbf{h}_{i}^{T}\mathbf{h}_{j})}{\sum_{k}\exp(\tau_{ik}\cdot\mathbf{h}_{i}^{T}\mathbf{h}_{k})}$, where the pairs suffering more from the collapse (i.e., $\mathbf{h}_{i}^{T}\mathbf{h}_{j}\rightarrow 1$) are balanced with a higher intensity. $\tau_{ij}$ is a temperature reflecting the relations between $i$ and $j$ estimated from the interaction data, and we propose two adaptive temperature designs to adapt to the pairs with varying degrees of correlations:
@@
    We use a unparameterrized polynomial graph filter $\phi(\cdot)=\sum_{l=0}^{L}\mathbf{\hat{A}}^{l}$ to measure the relation: $\tau_{uv}=\pi(\cos(\phi(u),\phi(i))$), where $\cos(\cdot,\cdot)$ is the cosine similarity. The user/item pairs with a higher similarity indicating a closer relation should be optimized with a smaller temperature (i.e., lower intensity). Thus, we use a linear function $\pi(\cdot)$ mapping the similarity to a proper range $[\tau_{0},\tau_{1}]$: the highest and lowest similarity is mapped to $\tau_{0}$ and $\tau_{1}$, respectively.
We will compare the two designs in Section 5.3.2. We substitute Equation (20) into Algorithm 1 and obtain the running algorithm for DirectSpec+.
4.4. Discussion
4.4.1. Connection with SSL-based Methods
Self-supervised learning (SSL) has shown great potential in recent years. Some SSL methods can generate high-quality representations without explicitly sampling negative samples that have been applied to recommendation as well, such as Barlow Twins (BT) (Zbontar et al., 2021), LogDet (Chen et al., 2023), and spectral contrastive loss (SCL) (HaoChen et al., 2021):
(21) $\mathcal{L}_{BT}=\sum_{i}\left(1-\mathcal{C}_{ii}\right)^{2}+\lambda\sum_{i}\sum_{j\neq i}\mathcal{C}_{ij}^{2},$
$\mathcal{L}_{logdet}=\operatorname{Tr}\left(\Sigma\right)-\log\det\left(\Sigma\right),$
$\mathcal{L}_{SCL}=\sum_{i}\sum_{i\neq j}\left(\mathbf{h}_{i}^{T}\mathbf{h}_{j}\right)^{2},$
where $\mathcal{C}_{ij}=\mathbf{H}[:,i]^{T}\mathbf{H}[:,j]$ is the covariance, and $\Sigma=\mathbf{H}^{T}\mathbf{H}$ is the covariance matrix.
Proposition 4.
BT, LogDet, and SCL are special cases of DirectSpec (Equation (11)) with $f\left(\sigma_{1},\cdots,\sigma_{d}\right)=\sum_{i}(\sigma_{i}^{2}-1)^{2}$ for BT and SCL and $f\left(\sigma_{1},\cdots,\sigma_{d}\right)=\sum_{i}\sigma_{i}-2\log\sigma_{i}$ for LogDet, respectively.
With $\mathbf{H}$ being normalized column-wisely, $\sum\sigma_{i}^{2}=\operatorname{Tr}(\mathbf{Q}diag(\sigma_{i}^{2})\mathbf{Q}^{T})=\operatorname{Tr}(\mathbf{H}^{T}\mathbf{H})=d$. Then, Barlow Twins can be rewritten as follows:
(22) $\mathcal{L}_{BT}$ $=\operatorname{Tr}\left(\left(\mathbf{H}^{T}\mathbf{H}-\mathbf{I}\right)^{2}\right)=\operatorname{Tr}\left(\mathbf{Q}diag\left(\sigma_{i}^{2}\right)\mathbf{Q}^{T}-2\mathbf{Q}diag\left(\sigma_{i}\right)\mathbf{Q}^{T}+\mathbf{Q}\mathbf{Q}^{T}\right)$
$=\operatorname{Tr}\left(\mathbf{Q}diag\left((\sigma_{i}^{2}-1)^{2}\right)\mathbf{Q}^{T}\right)=\sum_{i}\left(\sigma_{i}^{2}-1\right)^{2}.$
With $\mathbf{H}$ being normalized row-wisely, we can rewrite SCL as a similar form to BT:
(23) $\mathcal{L}_{SCL}=\operatorname{Tr}\left(\left(\mathbf{H}\mathbf{H}^{T}-\mathbf{I}\right)^{2}\right)=\operatorname{Tr}\left(\mathbf{P}diag\left((\sigma_{i}^{2}-1)^{2}\right)\mathbf{P}^{T}\right)=\sum_{i}\left(\sigma_{i}^{2}-1\right)^{2}.$
Similarly, we can reformulate LogDet as:
(24) $\mathcal{L}_{logdet}=\operatorname{Tr}\left(\mathbf{H}^{T}\mathbf{H}\right)-\log\det\left(\mathbf{H}^{T}\mathbf{H}\right)=\sum_{i}\sigma_{i}^{2}-\log\prod\sigma_{i}^{2}=\sum_{i}\sigma_{i}^{2}-2\log\sigma_{i}.$
By substitute Equation (22) (or (23)) and (24) into Equation (11) and solve them, we can get the minima achieved at $\sigma_{1}=\cdots=\sigma_{d}=\sqrt{d}$ for the three methods, $\beta=1-2d$ and $\beta=\frac{1-d}{d}$ for BT/SCL and LogDet when $c=d$, respectively, demonstrating that they can prevent the representations from collapsing by implicitly balancing the embedding spectrum. Since they belong to the DirectSpec framework, our proposed DirectSpec+ balances spectrum more adaptively and effectively, showing superiority over them. It is worth noting that InfoNCE loss cannot be reformulated as a case of DirectSpec. However, the embeddings are nearly and tend to be more orthogonal to each other as the dimensionality $d$ becomes larger, indicating a perfectly balanced spectrum, as shown in Proposition 3 of (Wang and Isola, 2020). We will compare our proposed methods with them in Section 5. Given DirectSpec’s generality and similarity to SSL methods that have been applied across various domains, we believe that DirectSpec can also contribute to other domains such as images, graph-structured data, etc.
Table 1. Comparison of time complexity.
Model DirectSpec Barlow Twins SCL LogDet
Complexity $c|\mathbf{R}^{+}|Bd$ $c|\mathbf{R}^{+}|d^{2}$ $c|\mathbf{R}^{+}|Bd$ $c|\mathbf{R}^{+}|d^{2}\left(1+\frac{d}{B}\right)$
Table 2. Statistics of datasets.
@@ 4.4.1. Connection with SSL-based Methods
4.4.2. Complexity
We compare the complexity of DirectSpec+ with Barlow Twins, spectral contrastive loss, and LetDet in Table 1, where $B$ and $c$ denote the batch size and the required epochs, respectively. We ignore the complexity for optimizing the alignment here. Overall, LogDet has the most expensive computational cost as it requires computing the matrix determinant with complexity as $\mathcal{O}(d^{3})$. The complexity of DirectSpec+ is identical to that of SCL. The difference of complexity between DirectSpec+ and BT lies in the term $Bd$ and $d^{2}$, where DirectSpec+ might be more costly when $B$ is large. However, considering that we set $B=256$ in the experiments, and the complexity of back propagation for BT and SCL is still $c|\mathbf{R}^{+}|Bd$ which is not contained in Table 1, DirectSpec+ is more time-efficient than BT and SCL.
4.4.3. Comparison between Embedding Collapse and Over-Smoothing
Generally, embedding collapse can be mathematically formulated as follows:
(25) ${\rm rank}\left(\mathbf{H}^{(L)}\right)<\epsilon,$
where $\mathbf{H}^{(L)}$ is the embedding when erank converges after updating $L$ times, $\epsilon<d$, and a smaller $\epsilon$ indicates more serious embedding collapse. Over-smoothing is attributed to the shrunk graph spectrum as repeating the message passing that can also be formulated as Equation (25). The collapse issue analyzed in this work shares similarities with over-smoothing, due to the observation that the parameter-updating process in recommendation algorithm shares the same form as the message passing in GNN. However, since the cause of over-smoothing in GNN lies in the graph spectrum, existing works tackle this issue mostly by dealing with the graph spectrum, such as randomly dropping out the edges to perturb the spectrum (Rong et al., 2020), balancing the contribution of different order neighborhood (Gasteiger et al., 2019; Peng et al., 2022a) to avoid all eigenvalues from shrinking, etc., which fail to solve the embedding collapse in this work, as it exists generally on recommendation algorithms without relying on explicit graph structures. On the other hand, our proposed DirectSpec tackles a general embedding collapse by directly balancing the embedding spectrum which is proved to solve this issue. We will show that our DirectSpec is also effective on over-smoothing and embedding collapse caused by other reasons in Section 5.
4.4.4. Comparison with Related Works
Recently, several efforts have been made to address ‘collapse issues’ in recommendation methods (Chen et al., 2023; Zhang et al., 2023; Guo et al., 2024), which shares similarity with our work. Guo et al. (Guo et al., 2023) points out an embedding collapse within the feature interaction module for click-rate prediction (CTR) caused by the imbalanced gradients over different singular vectors, while the embedding collapse issue in our work stems from optimizing a CF algorithm. The dimensional collapse in graph collaborative filtering (GCF) investigated in (Chen et al., 2023; Zhang et al., 2023) overlap more significantly with ours. Both the works exploit a LogDet function to address the issue, which are more computationally expensive than DirectSpec as analyzed in Section 4.4.2. We will demonstrate the superiority of our proposed methods in Section 5.
5. Experiments
In this section, we comprehensively evaluate our proposed methods on three public datasets. We implement DirectSpec and DirectSpec+ on two popular baselines: MF and LightGCN. Particularly, we introduce data description and implementation details in Section 5.1; we compare the proposed methods with other competitive baselines in terms of accuracy and efficiency in Section 5.2. Finally, we conduct model analysis in Section 5.3, including detailed analysis and experimental results on how DirectSpec+ prevents embeddings from collapsing without relying on negative samples, and how different settings of hyperparameters affect model performance.
5.1. Experimental Settings
5.1.1. Datasets